two squares have the perimeters of 20 inches and 48 inches. How much longer is a side of the larger square than a side of the smaller squre. plases hlep
Answer by josgarithmetic(39617)

perimeter 20.
x for length of side
4x=20
x=5

perimeter 48.
y for length of side
4y=48
y=12

How much longer?
y-x=12-5=7